Different cash flow. Given the following cash inflow, what is the present value of this cash flow at 8%, 14%, and 27% discount rates? Year 1 Year 2: Years 3 through 7: Year 8: $1,000 $5,000 $0 $25,000

A) What is the present value of this cash flow at 8​% discount​ rate?

B) What is the present value of this cash flow at 14​% discount​ rate?

C) What is the present value of this cash flow at 27​% discount​ rate?
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

a.Present value=1000/1.08+5000/1.08^2+25000/1.08^8

=$18719.34(Approx).

b.Present value=1000/1.14+5000/1.14^2+25000/1.14^8

=$13488.51(Approx).

c.Present value=1000/1.27+5000/1.27^2+25000/1.27^8

=$7581.52(Approx).